OpenStudy (anonymous):

Two cars that are 150 miles apart start driving toward each other on parallel roads. The average speed of the first car is 60 miles per hour. The average speed of the second car is 55 miles per hour. Which equation can be used to determine t, the time it takes for the two cars to pass each other?

OpenStudy (anonymous):

a. 60t – 55t = 0 b. 60t + 55t = 1 c. 60t + 55t = 150 d. 60t – 55t = 150

OpenStudy (anonymous):

i think c

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

yah it's c

OpenStudy (shinalcantara):

remember that velocity is distance over time or the distance is velocity times time since the total distance is 150 then you'll have it C